### v1

# Objective and subjective evaluation of speech enhancement methods in the UDASE task of the 7th CHiME challenge

Abstract
Supervised models for speech enhancement are trained using artificially generated mixtures of clean speech and noise signals. However, the synthetic training conditions may not accurately reflect real-world conditions encountered during testing. This discrepancy can result in poor performance when the test domain significantly differs from the synthetic training domain. To tackle this issue, the UDASE task of the 7th CHiME challenge aimed to leverage real-world noisy speech recordings from the test domain for unsupervised domain adaptation of speech enhancement models. Specifically, this test domain corresponds to the CHiME-5 dataset, characterized by real multi-speaker and conversational speech recordings made in noisy and reverberant domestic environments, for which ground-truth clean speech signals are not available. In this paper, we present the objective and subjective evaluations of the systems that were submitted to the CHiME-7 UDASE task, and we provide an analysis of the results. This analysis reveals a limited correlation between subjective ratings and several supervised nonintrusive performance metrics recently proposed for speech enhancement. Conversely, the results suggest that more traditional intrusive objective metrics can be used for in-domain performance evaluation using the reverberant LibriCHiME-5 dataset developed for the challenge. The subjective evaluation indicates that all systems successfully reduced the background noise, but always at the expense of increased distortion. Out of the four speech enhancement methods evaluated subjectively, only one demonstrated an improvement in overall quality compared to the unprocessed noisy speech, highlighting the difficulty of the task. The tools and audio material created for the CHiME-7 UDASE task are shared with the community.
keywords:
CHiME challenge, multi-speaker conversational speech, speech enhancement, unsupervised domain adaptation, ITU-T P.835 listening test.
[…]
1 Introduction
The speech enhancement task consists of improving the quality and intelligibility of a degraded speech signal recording. One approach to achieve this is through noise suppression algorithms, which aim to estimate the clean speech signal by removing the additive background noise in the recording. Over the past 50 years, traditional signal-processing-based speech enhancement algorithms (Boll, 1979; Lim & Oppenheim, 1979; Ephraim & Malah, 1984; Martin, 2005; Loizou, 2013) have been progressively outperformed by data-driven approaches using hidden Markov models (HMMs) (Sameti et al., 1998; Ephraim, 1992; Sameti et al., 1998), codebook-based approaches (Srinivasan et al., 2005), non-negative matrix factorization (NMF) (Mohammadiha et al., 2013), and more recently deep learning (Wang & Chen, 2018).
There has been great progress in speech enhancement in recent years thanks to deep learning models trained in a supervised manner. The conventional approach in supervised speech enhancement involves three main ingredients:
  1. 1.
    A model, which provides a prediction of the clean speech signal given the noisy recording. Current state-of-the-art methods rely on deep neural networks.
  2. 2.
    A metric, which measures the discrepancy between the clean speech estimate and the ground-truth signal. During training, the metric corresponds to a differentiable loss function, which is minimized to estimate the model parameters. At test time, the metric (which can differ from the training loss function) is used to evaluate the performance of the trained model.
  3. 3.
    A labeled dataset, which consists of noisy speech signals paired with their corresponding clean reference signals. During training, these clean reference signals serve as the model’s targets, and at test time, they are used to compute intrusive performance metrics.
Unfortunately, it is very difficult, if not impossible, to acquire labeled noisy speech signals in real-world conditions due to cross-talk between microphones. Datasets for supervised learning and performance evaluation with intrusive metrics have to be generated artificially, by creating synthetic mixtures of isolated speech and noise signals.
A large research effort has been put recently on the three main ingredients of supervised speech enhancement (the model, the metric, and the labeled dataset), and it is undeniable that this effort has led to unprecedented results, e.g., Weninger et al. (2015); Fu et al. (2017); Pascual et al. (2017); Choi et al. (2018); Zhao et al. (2018); Fu et al. (2019); Défossez et al. (2020); Cosentino et al. (2020); Hao et al. (2021); Pandey & Wang (2021); Fu et al. (2021); Richter et al. (2023). However, the fully supervised learning paradigm also has limitations when applied to speech enhancement. First, creating a synthetic dataset of realistic noisy speech mixtures is not easy and requires important engineering efforts. Second, supervised speech enhancement is effective as long as the acoustic recording conditions at test time are covered by the synthetic training data. This is a constraint that is hard to satisfy due to the variability of the acoustic recording conditions, in terms of noise type, signal-to-noise ratio, recording equipment, speaker-to-microphone distance and orientation, reverberation, etc. Supervised speech enhancement performance can thus decrease significantly in case of mismatch between the training and testing conditions (Pandey & Wang, 2020; Bie et al., 2022; Richter et al., 2023; Gonzalez et al., 2023). Finally, when the test domain differs from the synthetic training domain, supervised learning necessitates rebuilding the synthetic training dataset and retraining the model, which is time-consuming and computationally intensive. A more effective approach would be to automatically adapt models to real, unlabeled noisy speech recordings, eliminating the need for ground-truth clean speech signals.
The ability to adapt to unknown adverse acoustic conditions while perceiving speech is a fundamental property of the human auditory system (Bent et al., 2009; Brandewie & Zahorik, 2010; Cooke et al., 2022), which cannot be reproduced in machine listening systems using a fully supervised approach. Adaptation of speech enhancement systems using real unlabeled noisy speech data is the main challenge that the CHiME-7 UDASE [^1] Unsupervised Domain Adaptation for conversational Speech Enhancement. task tried to address (Leglaive et al., 2023). Previous challenges for single-channel speech enhancement, such as the deep noise suppression (DNS) challenges (Reddy et al., 2020, 2021a, 2021b; Dubey et al., 2022), focused on a supervised setup with a training set consisting of a large amount of labeled synthetic data intended to cover diverse conditions. The CHiME-7 UDASE task was intended to study a different situation, targeting single-channel speech enhancement in a specific domain for which no well-matched labeled data are available for training.
In Leglaive et al. (2023), we introduced the task and described the data and the baseline system. In the present paper, we extend the description of the CHiME-7 UDASE task by introducing the speech enhancement methods that were submitted to the challenge, describing their objective and subjective evaluation, and providing an analysis of the results. Along with this paper, we release the Javascript experimental platform we developed for the listening test, following the ITU-T P.835 (2003) methodology. We also release the audio files that were submitted to the challenge along with the corresponding human listening scores (in addition to various objective evaluation scores), which could serve as a voice quality dataset for speech enhancement research. [^2] https://doi.org/10.5281/zenodo.10418311 To the best of our knowledge, this is the first dataset of ITU-T P.835 (2003) subjective evaluation results made publicly available.
The paper is organized as follows. The task and datasets are presented in Section 2. The objective and subjective evaluation protocols are described in Sections 3 and 4, respectively. Section 5 introduces the speech enhancement methods that participated in the CHiME-7 UDASE task. The results of the subjective and objective evaluations are presented and analyzed in Section 6 before concluding in Section 7.
2 Task and datasets
2.1 Task
The CHiME-7 UDASE task focuses on single-channel speech enhancement in a specific target domain for which no well-matched labeled training data are available. It consists of using unlabeled data in the target domain to adapt supervised speech enhancement models trained on synthetic labeled data in a source domain. The target domain corresponds to the real conversational speech recordings of the CHiME-5 dataset (Barker et al., 2018). These recordings were made during dinner parties in real homes, with multiple speakers in noisy and reverberant environments. Given a mixture of one or more reverberant speakers and additive background noise, the objective of this task is to estimate the clean, potentially multi-speaker, reverberant speech, removing the additive background noise. The task is motivated by an assistive listening use case, in which a speech enhancement algorithm can help any individual to better engage in a conversation, by improving the overall speech quality and intelligibility within the ambient noise.
2.2 Datasets
The CHiME-7 UDASE task is built upon three datasets:
  1. 1.
    The CHiME-5 dataset (Barker et al., 2018) for the in-domain unlabeled data, which are used for training, development, and evaluation;
  2. 2.
    The LibriMix dataset (Cosentino et al., 2020) for the out-of-domain labeled data, which are used for training and development only;
  3. 3.
    A new reverberant LibriCHiME-5 dataset, which was created to provide “close to in-domain” labeled data for development and evaluation only.
All three datasets include mixtures of reverberant speech and noise, with up to three overlapping speakers. All signals are sampled at 16 kHz. The datasets are briefly presented in the rest of this section, and additional information can be found in Leglaive et al. (2023).
2.3 CHiME-5 in-domain unlabeled data
The CHiME-5 dataset (Barker et al., 2018) consists of recordings made during twenty real dinner parties (or sessions) of between two and three hours. Each dinner party involved four participants wearing binaural microphones and took place in a different home, with three recording locations per home (kitchen, dining room, living room). The CHiME-5 recordings include natural conversations between multiple speakers in reverberant and noisy environments, and they are fully transcribed. Using the CHiME-5 transcription files, we estimated that 22% of the audio recordings contain only noise, 51% contain one single active speaker, and 20%, 5%, and 2% contain two, three, and four overlapping speakers, respectively. [^3] These numbers were obtained without considering any constraint on the overlap duration and using a labeling precision of 0.01 second. The CHiME-7 UDASE task only uses the binaural recordings (Kinect recordings are also available), from which single-channel audio segments with up to 3 overlapping speakers and background noise were extracted.
The dataset contains training, development, and evaluation sets, with different speakers in each set. The training set consists of the raw single-channel audio segments extracted from the binaural recordings when the participant wearing the microphone does not speak. It is intended to be used for unsupervised adaptation on in-domain unlabeled noisy speech data.
No ground-truth clean speech signals are available for the CHiME-5 noisy speech recordings. This is a major difficulty because for developing and evaluating a speech enhancement system one needs to compute objective performance metrics, which is usually achieved using ground-truth signals. To circumvent this difficulty, the transcription files were used to segment the CHiME-5 recordings in short audio segments labeled with the maximum number of simultaneously active speakers (0, 1, 2 or 3). [^4] This only corresponds to a maximum value, i.e., through the duration of a segment the number of simultaneously-active speakers can vary between 0 and the maximum value. Moreover, a segment might contain more speakers than the labeled maximum number of simultaneously active speakers. For instance, a segment labeled as a single-speaker might contain two active speakers who do not speak simultaneously. This segmentation was done only for the development and evaluation sets, which simulates the reasonable scenario where we can afford to manually annotate a small amount of data with speaker count labels for development and evaluation, but this procedure cannot be easily done for a large training set. The noise-only segments were subsequently used to create the synthetic labeled noisy speech mixtures of the reverberant LibriCHiME-5 dataset (see Section 2.5). The single-speaker segments can be used to compute nonintrusive (reference-free) objective performance metrics, meaning they do not require ground-truth clean speech signals. The dataset also contains an evaluation subset intended to be used for the listening test. It consists of test audio samples extracted by looking for segments of 4 to 5 seconds with at least 3 seconds of speech and 0.25 seconds without speech at the beginning and the end. Additional constraints were taken into account to ensure a balanced subset in terms of the speaker’s gender, recording location, and session.
The segmented CHiME-5 dataset used for the CHiME-7 UDASE task is summarized in Table 2.3. Additional details about the segmentation of the CHiME-5 original recordings are provided in Leglaive et al. (2023), and the scripts to generate the dataset are available online. [^5] https://github.com/UDASE-CHiME2023/CHiME-5
Table 1: Segmented CHiME-5 dataset. Dev and eval subsets are labeled with the maximum number of simultaneously active speakers (0, 1, 2, 3). eval/LT corresponds to the evaluation subset for the listening test.
Figure 2: Instructions and rating scales for the ITU-T P.835 listening test.
8pt. Sample length (s) Total duration
Subset # samples Mean STD ($H$:$M$:$S$)
train 27 517 10.91 14.10 83:22:29
dev/0 912 6.50 4.10 1:38:49
dev/1 5 719 5.89 3.49 9:21:53
dev/2 3 835 5.23 2.43 5:34:33
dev/3 667 4.61 1.84 0:51:14
eval/0 977 5.73 3.35 1:33:19
eval/1 3 013 5.54 2.94 4:35:05
eval/2 1 552 4.88 2.04 2:06:07
eval/3 233 4.21 1.17 0:16:21
eval/LT 241 4.72 0.34 0:18:58
8pt.
The design of the listening test for the CHiME-7 UDASE task followed the ITU-T P.835 (2003) methodology (Loizou, 2011). We were inspired by Hu & Loizou (2006, 2007); Naderi & Cutler (2021) who also evaluated speech enhancement algorithms following this methodology.
Table 1: Segmented CHiME-5 dataset. Dev and eval subsets are labeled with the maximum number of simultaneously active speakers (0, 1, 2, 3). eval/LT corresponds to the evaluation subset for the listening test.
Figure 2: Instructions and rating scales for the ITU-T P.835 listening test.
Figure 5: Boxplots, violin plots, and mean results for the subjective BAK (4(a)), SIG (4(b)), and OVRL (4(c)) mean opinion scores of the ITU-T P.835 listening test. Black dots and numbers above the box/violin plots correspond to the mean results. In each figure, the systems are ranked according to their mean results.