### v2

# Objective and subjective evaluation of speech enhancement methods in the UDASE task of the 7th CHiME challenge

Abstract
Supervised models for speech enhancement are trained using artificially generated mixtures of clean speech and noise signals. However, the synthetic training conditions may not accurately reflect real-world conditions encountered during testing. This discrepancy can result in poor performance when the test domain significantly differs from the synthetic training domain. To tackle this issue, the UDASE task of the 7th CHiME challenge aimed to leverage real-world noisy speech recordings from the test domain for unsupervised domain adaptation of speech enhancement models. Specifically, this test domain corresponds to the CHiME-5 dataset, characterized by real multi-speaker and conversational speech recordings made in noisy and reverberant domestic environments, for which ground-truth clean speech signals are not available. In this paper, we present the objective and subjective evaluations of the systems that were submitted to the CHiME-7 UDASE task, and we provide an analysis of the results. This analysis reveals a limited correlation between subjective ratings and several supervised nonintrusive performance metrics recently proposed for speech enhancement. Conversely, the results suggest that more traditional intrusive objective metrics can be used for in-domain performance evaluation using the reverberant LibriCHiME-5 dataset developed for the challenge. The subjective evaluation indicates that all systems successfully reduced the background noise, but always at the expense of increased distortion. Out of the four speech enhancement methods evaluated subjectively, only one demonstrated an improvement in overall quality compared to the unprocessed noisy speech, highlighting the difficulty of the task. The tools and audio material created for the CHiME-7 UDASE task are shared with the community.
keywords:
CHiME challenge, multi-speaker conversational speech, speech enhancement, unsupervised domain adaptation, ITU-T P.835 listening test.
[…]
1 Introduction
The speech enhancement task consists of improving the quality and intelligibility of a degraded speech signal recording. One approach to achieve this is through noise suppression algorithms, which aim to estimate the clean speech signal by removing the additive background noise in the recording. Over the past 50 years, traditional signal-processing-based speech enhancement algorithms [Boll, 1979, Lim & Oppenheim, 1979, Ephraim & Malah, 1984, Martin, 2005, Loizou, 2013] have been progressively outperformed by data-driven approaches using hidden Markov models (HMMs) [Sameti et al., 1998, Ephraim, 1992, Sameti et al., 1998], codebook-based approaches [Srinivasan et al., 2005], non-negative matrix factorization (NMF) [Mohammadiha et al., 2013], and more recently deep learning [Wang & Chen, 2018].
There has been great progress in speech enhancement in recent years thanks to deep learning models trained in a supervised manner. The conventional approach in supervised speech enhancement involves three main ingredients:
  1. 1.
    A model, which provides a prediction of the clean speech signal given the noisy recording. Current state-of-the-art methods rely on deep neural networks.
  2. 2.
    A metric, which measures the discrepancy between the clean speech estimate and the ground-truth signal. During training, the metric corresponds to a differentiable loss function, which is minimized to estimate the model parameters. At test time, the metric (which can differ from the training loss function) is used to evaluate the performance of the trained model.
  3. 3.
    A labeled dataset, which consists of noisy speech signals paired with their corresponding clean reference signals. During training, these clean reference signals serve as the model’s targets, and at test time, they are used to compute intrusive performance metrics.
Unfortunately, it is very difficult, if not impossible, to acquire labeled noisy speech signals in real-world conditions due to cross-talk between microphones. Datasets for supervised learning and performance evaluation with intrusive metrics have to be generated artificially, by creating synthetic mixtures of isolated speech and noise signals.
A large research effort has been put recently on the three main ingredients of supervised speech enhancement (the model, the metric, and the labeled dataset), and it is undeniable that this effort has led to unprecedented results, e.g., Weninger et al. [2015], Fu et al. [2017], Pascual et al. [2017], Choi et al. [2018], Zhao et al. [2018], Fu et al. [2019], Défossez et al. [2020], Cosentino et al. [2020], Hao et al. [2021], Pandey & Wang [2021], Fu et al. [2021], Richter et al. [2023]. However, the fully supervised learning paradigm also has limitations when applied to speech enhancement. First, creating a synthetic dataset of realistic noisy speech mixtures is not easy and requires important engineering efforts. Second, supervised speech enhancement is effective as long as the acoustic recording conditions at test time are covered by the synthetic training data. This is a constraint that is hard to satisfy due to the variability of the acoustic recording conditions, in terms of noise type, signal-to-noise ratio, recording equipment, speaker-to-microphone distance and orientation, reverberation, etc. Supervised speech enhancement performance can thus decrease significantly in case of mismatch between the training and testing conditions [Pandey & Wang, 2020, Bie et al., 2022, Richter et al., 2023, Gonzalez et al., 2023]. Finally, when the test domain differs from the synthetic training domain, supervised learning necessitates rebuilding the synthetic training dataset and retraining the model, which is time-consuming and computationally intensive. A more effective approach would be to automatically adapt models to real, unlabeled noisy speech recordings, eliminating the need for ground-truth clean speech signals.
The ability to adapt to unknown adverse acoustic conditions while perceiving speech is a fundamental property of the human auditory system [Bent et al., 2009, Brandewie & Zahorik, 2010, Cooke et al., 2022], which cannot be reproduced in machine listening systems using a fully supervised approach. Adaptation of speech enhancement systems using real unlabeled noisy speech data is the main challenge that the CHiME-7 UDASE [^1] Unsupervised Domain Adaptation for conversational Speech Enhancement. task tried to address [Leglaive et al., 2023]. Previous challenges for single-channel speech enhancement, such as the deep noise suppression (DNS) challenges [Reddy et al., 2020, 2021a, 2021b, Dubey et al., 2022], focused on a supervised setup with a training set consisting of a large amount of labeled synthetic data intended to cover diverse conditions. The CHiME-7 UDASE task was intended to study a different situation, targeting single-channel speech enhancement in a specific domain for which no well-matched labeled data are available for training.
In Leglaive et al. [2023], we introduced the task and described the data and the baseline system. In the present paper, we extend the description of the CHiME-7 UDASE task by introducing the speech enhancement methods that were submitted to the challenge, describing their objective and subjective evaluation, and providing an analysis of the results. Along with this paper, we release the JavaScript experimental platform we developed for the listening test, following the ITU-T P.835 [2003] methodology. We also release the audio files that were submitted to the challenge along with the corresponding human listening scores (in addition to various objective evaluation scores), which could serve as a voice quality dataset for speech enhancement research [Leglaive et al., 2024]. To the best of our knowledge, this is the first dataset of ITU-T P.835 [2003] subjective evaluation results made publicly available.
The paper is organized as follows. The task and datasets are presented in Section 2. The objective and subjective evaluation protocols are described in Sections LABEL:sec:objective_eval and LABEL:sec:listening_test, respectively. Section LABEL:sec:SE_methods introduces the speech enhancement methods that participated in the CHiME-7 UDASE task. The results of the subjective and objective evaluations are presented and analyzed in Section LABEL:sec:results before concluding in Section 7.
2 Task and datasets
2.1 Task
The CHiME-7 UDASE task focuses on single-channel speech enhancement in a specific target domain for which no well-matched labeled training data are available. It consists of using unlabeled data in the target domain to adapt supervised speech enhancement models trained on synthetic labeled data in a source domain. The target domain corresponds to the real conversational speech recordings of the CHiME-5 dataset [Barker et al., 2018]. These recordings were made during dinner parties in real homes, with multiple speakers in noisy and reverberant environments. Given a mixture of one or more reverberant speakers and additive background noise, the objective of this task is to estimate the clean, potentially multi-speaker, reverberant speech, removing the additive background noise. The task is motivated by an assistive listening use case, in which a speech enhancement algorithm can help any individual to better engage in a conversation, by improving the overall speech quality and intelligibility within the ambient noise.
2.2 Datasets
The CHiME-7 UDASE task is built upon three datasets:
  1. 1.
    The CHiME-5 dataset [Barker et al., 2018] for the in-domain unlabeled data, which are used for training, development, and evaluation;
  2. 2.
    The LibriMix dataset [Cosentino et al., 2020] for the out-of-domain labeled data, which are used for training and development only;
  3. 3.
    A new reverberant LibriCHiME-5 dataset, which was created to provide “close to in-domain” labeled data for development and evaluation only.
All three datasets include mixtures of reverberant speech and noise, with up to three overlapping speakers. All signals are sampled at 16 kHz. The datasets are presented in the rest of this section, and additional information can be found in Leglaive et al. [2023].
2.3 CHiME-5 in-domain unlabeled data
The CHiME-5 dataset [Barker et al., 2018] consists of recordings made during twenty real dinner parties (or sessions) of between two and three hours. Each dinner party involved four participants wearing binaural microphones and took place in a different home, with three recording locations per home (kitchen, dining room, living room). The CHiME-5 recordings include natural conversations between multiple speakers in reverberant and noisy environments, and they are fully transcribed. Using the CHiME-5 transcription files, we estimated that 22% of the audio recordings contain only noise, 51% contain one single active speaker, and 20%, 5%, and 2% contain two, three, and four overlapping speakers, respectively. [^2] These numbers were obtained without considering any constraint on the overlap duration and using a labeling precision of 0.01 second. The CHiME-7 UDASE task only uses the binaural recordings (Kinect recordings are also available), from which single-channel audio segments with up to 3 overlapping speakers and background noise were extracted.
The dataset contains training, development, and evaluation sets, with different speakers in each set. The training set consists of the raw single-channel audio segments extracted from the binaural recordings when the participant wearing the microphone does not speak. It is intended to be used for unsupervised adaptation on in-domain unlabeled noisy speech data.
No ground-truth clean speech signals are available for the CHiME-5 noisy speech recordings. This is a major difficulty because for developing and evaluating a speech enhancement system one needs to compute objective performance metrics, which is usually achieved using ground-truth signals. To circumvent this difficulty, the transcription files were used to segment the CHiME-5 recordings in short audio segments labeled with the maximum number of simultaneously active speakers (0, 1, 2 or 3). [^3] This only corresponds to a maximum value, i.e., through the duration of a segment the number of simultaneously-active speakers can vary between 0 and the maximum value. Moreover, a segment might contain more speakers than the labeled maximum number of simultaneously active speakers. For instance, a segment labeled as a single-speaker might contain two active speakers who do not speak simultaneously. This segmentation was done only for the development and evaluation sets, which simulates the reasonable scenario where we can afford to manually annotate a small amount of data with speaker count labels for development and evaluation, but this procedure cannot be easily done for a large training set. The noise-only segments were subsequently used to create the synthetic labeled noisy speech mixtures of the reverberant LibriCHiME-5 dataset (see Section 2.5). The single-speaker segments can be used to compute nonintrusive (reference-free) objective performance metrics, meaning they do not require ground-truth clean speech signals. The dataset also contains an evaluation subset intended to be used for the listening test. It consists of test audio samples extracted by looking for segments of 4 to 5 seconds with at least 3 seconds of speech and 0.25 seconds without speech at the beginning and the end. Additional constraints were taken into account to ensure a balanced subset in terms of the speaker’s gender, recording location, and session.
The segmented CHiME-5 dataset used for the CHiME-7 UDASE task is summarized in Table 2.3. Additional details about the segmentation of the CHiME-5 original recordings are provided in Leglaive et al. [2023], and the scripts to generate the dataset are available online. [^4] https://github.com/UDASE-CHiME2023/CHiME-5 (last accessed: 2024-02-02)
8pt. Sample length (s) Total duration
Subset # samples Mean SD (HH:MM:SS)
train 27 517 10.91 14.10 83:22:29
dev/0 912 6.50 4.10 1:38:49
dev/1 5 719 5.89 3.49 9:21:53
dev/2 3 835 5.23 2.43 5:34:33
dev/3 667 4.61 1.84 0:51:14
eval/0 977 5.73 3.35 1:33:19
eval/1 3 013 5.54 2.94 4:35:05
eval/2 1 552 4.88 2.04 2:06:07
eval/3 233 4.21 1.17 0:16:21
eval/LT 241 4.72 0.34 0:18:58
8pt.
Table 1: Segmented CHiME-5 dataset. Dev and eval subsets are labeled with the maximum number of simultaneously active speakers (0, 1, 2, 3). eval/LT corresponds to the evaluation subset for the listening test.
2.4 LibriMix out-of-domain labeled data
The CHiME-7 UDASE task employs the LibriMix dataset [Cosentino et al., 2020] for supervised learning on out-of-domain data. LibriMix was originally developed for speech separation in noisy environments, and it is derived from LibriSpeech “clean” utterances [Panayotov et al., 2015] and WHAM! noises [Wichern et al., 2019]. The Libri2Mix and Libri3Mix versions of the dataset contain noisy speech mixtures with 2 and 3 overlapping speakers, respectively. A single-speaker version of LibriMix (Libri1Mix) can be obtained by simply discarding one of the two speakers in Libri2Mix mixtures. A complete description of LibriMix is provided by Cosentino et al. [2020].
2.5 Reverberant LibriCHiME-5 close-to-in-domain labeled data
8pt. RT60 (s)
Subset Home Room # RIRs Mean SD
dev 2 Living room 160 0.582 0.023
2 Kitchen 120 0.500 0.022
2 Bedroom 136 0.471 0.056
3 Living room 136 0.437 0.026
3 Bedroom 120 0.431 0.036
eval 3 Kitchen 128 0.451 0.037
4 Living room 144 0.460 0.034
4 Kitchen 128 0.375 0.035
4 Bedroom 128 0.387 0.022
8pt.
Table 2: Estimated RT60s (in seconds) of the rooms in the VoiceHome corpus [Bertin et al., 2016] that were used to create the reverberant LibriCHiME-5 data (dev and eval subsets). The column ‘# RIRs’ gives the number of single-channel RIRs used to compute the mean and standard deviation (SD) of the RT60.
We created the reverberant LibriCHiME-5 dataset to provide “close-to-in-domain” labeled data for computing standard intrusive objective performance metrics used in speech enhancement, such as the scale-invariant signal-to-distortion ratio (SI-SDR) [Le Roux et al., 2019]. This dataset consists of synthetic mixtures of reverberant speech and noise, with up to three simultaneously active speakers, labeled with the clean reference speech signals. In-domain noise signals were extracted from the CHiME-5 recordings using the ground-truth transcriptions (see Section 2.3), and clean speech utterances were taken from the LibriSpeech dataset [Panayotov et al., 2015] and convolved with room impulse responses (RIRs) from the VoiceHome corpus [Bertin et al., 2016]. These RIRs were recorded in 12 different rooms of 3 real homes, with 4 rooms per home: living room (room 1), kitchen (room 2), bedroom (room 3), and bathroom (room 4). Bathrooms were excluded for the reverberant LibriCHiME-5 dataset. In each room, RIRs were recorded for 2 different positions and geometries of an 8-channel microphone array and 7 to 9 different positions of the loudspeaker. Table 2.5 indicates the estimated reverberation times (RT60s) of the rooms in the VoiceHome corpus whose RIRs were used to create the reverberant LibriCHiME-5 dataset. The RT60 is defined as the time it takes for the sound energy to decrease by 60 dB after the extinction of the source. For each room in each home, we estimated the RT60 on each single-channel RIR by fitting a straight line on Schroeder’s integrated energy decay curve using linear least-squares regression [Schroeder, 1965]. We then computed the mean RT60 and the standard deviation for each room.
The process to create the synthetic reverberant LibriCHiME-5 dataset was the following. For each mixture in the dataset, we randomly chose the maximum number $n\in\{1,2,3\}$ of simultaneously active speakers in the mixture, with $p(n=i)=0.60,0.35,0.05$ for $i=1,2,3$, respectively, which is consistent with the distribution of the segmented CHiME-5 dataset. Each mixture’s speakers were randomly sampled from the list of LibriSpeech speakers with equal probability to be a male or a female. We used the VoiceHome corpus to simulate the acoustic recording environment in the reverberant LibriCHiME-5 dataset. For each mixture, we randomly and successively sampled a home, a room, an array position/geometry, $n$ speaker positions without replacement, and a channel of the microphone array, which gave a set of RIRs. LibriSpeech utterances were convolved with the selected RIRs to obtain the reverberant speech utterances. These were then mixed following speech activity patterns extracted from the CHiME-5 transcription files (diarization labels) to simulate a natural conversation between multiple speakers. This synthetic mixing of multi-speaker speech involved selecting and trimming LibriSpeech utterances in order to make them fit in the CHiME-5 activity patterns. The multi-speaker reverberant speech and noise mixtures were finally created such that the per-speaker signal-to-noise ratio (SNR) was distributed as a Gaussian with a mean of 5 dB and a standard deviation of 7 dB to match the SNR distribution of the CHiME-5 dataset as estimated by Brouhaha [Lavechin et al., 2023]. This was achieved by first sampling a global per-mixture SNR $x\sim\mathcal{N}(5,\sigma_{1}^{2})$ and then sampling a local per-speaker SNR $y_{n}\sim\mathcal{N}(x,\sigma_{2}^{2})$, with $\sigma_{1}=6.7082$ and $\sigma_{2}=2$ ($\sqrt{\sigma_{1}^{2}+\sigma_{2}^{2}}\approx 7$ dB). The value of $\sigma_{2}$ was chosen such that the loudness difference between multiple speakers remained moderate; this was again to simulate a conversation. A detailed description of the process implemented to create the reverberant LibriCHiME-5 mixtures (metadata and audio files) and the corresponding Python code are available online. [^5] https://github.com/UDASE-CHiME2023/reverberant-LibriCHiME-5 (last accessed: 2024-04-03).
Overall, in the reverberant LibriCHiME-5 dataset, the speech utterances were convolved with RIRs measured in real homes, the noise signals were extracted from the in-domain CHiME-5 recordings, the SNR was chosen to approximately match that of the target domain, and the speech utterances were mixed to simulate a conversation using the CHiME-5 transcription. It is therefore hoped that the performance on the reverberant LibriCHiME-5 dataset corresponds to an estimate of the performance on the CHiME-5 dataset, which will be confirmed by the analysis of the evaluation results. The reverberant LibriCHiME-5 dataset is summarized in Table LABEL:tab:librichime-5.
(b) SIG MOS results.
(c) OVRL MOS results.
Figure 4: Boxplots, violin plots, and mean results for the subjective BAK (LABEL:fig:listening_test_BAK), SIG (4(b)), and OVRL (4(c)) mean opinion scores of the ITU-T P.835 listening test. Black dots and numbers above the box/violin plots correspond to the mean results. In each figure, the systems are ranked according to their mean results.
6.2.5 Comparison with the objective evaluation
Figure 5: Pearson correlation between objective and subjective evaluation results, computed using only single-speaker samples (115 samples in total).
It can be interesting to compare the results of the objective and subjective evaluations. The CMGAN submissions, which were found to perform best in terms of DNSMOS metrics on both CHiME-5 and reverberant LibriCHiME-5 datasets, are the worst-performing systems according to the listening test (see Figure 2.3). As detailed in Close et al. [2023], this is probably due to the training of the CMGAN models that rely on the optimization of the supervised DNSMOS metrics. It appears that the models learned to maximize the metrics while providing output speech signals of unsatisfactory quality. This is probably because the DNSMOS model is supervised and it can thus provide unexpected or incoherent results when fed with signals that are too far from its training conditions. On the contrary, the objective evaluation results using standard objective performance metrics (SI-SDR, PESQ, and STOI) computed on the close-to-in-domain reverberant LibriCHiME-5 dataset are consistent with the subjective evaluation results on the CHiME-5 dataset, in terms of ranking of the methods. This again confirms that reverberant LibriCHiME-5 can be used to approximate in-domain performance in the context of the CHiME-7 UDASE task.
Finally, Figure 5 shows the PCC between the subjective SIG, BAK, OVRL MOS, and the nonintrusive DNSMOS and TorchAudio-Squim metrics. It can be seen that, apart from a PCC of 0.73 between the DNSMOS BAK score and the subjective BAK MOS, the nonintrusive objective metrics poorly correlate with the subjective ones (PCC lower than or equal to 0.50). The DNSMOS BAK score was also found to be the most reliable nonintrusive metric when analyzing the objective evaluation results in Section LABEL:subsec:obj_eval. Interestingly, the correlation between the subjective SIG, BAK, and OVRL metrics (0.17, 0.76, 0.66) is very close to the correlation between the equivalent objective metrics as provided by DNSMOS (0.18, 0.69, 0.66). This suggests that DNSMOS has appropriately learned the correlation between the SIG, BAK, and OVRL scores, but unfortunately, this is not sufficient to ensure good generalization, in particular on the CHiME-7 UDASE data.
7 Conclusion
Fully-supervised speech enhancement models are trained – and most of the time also evaluated – using only synthetic data, which cannot always capture the diversity of real-world speech recordings. A strong mismatch between the synthetic training domain and the real test domain will significantly affect the performance of the model. To address this issue, the CHiME-7 UDASE task aimed to leverage real-world unlabeled noisy speech recordings from the test domain for the unsupervised adaptation of speech enhancement models. Evaluating unsupervised domain adaptation methods for speech enhancement is by definition a challenging task because the ground-truth clean speech signals in the target domain are not available. In this paper, we presented the methodology and analyzed the results of the objective and subjective evaluations conducted in the CHiME-7 UDASE task.
Objective evaluation in the target domain (defined by the CHiME-5 recordings) relied on the recent DNSMOS P.835 and TorchAudio-Squim nonintrusive performance metrics. Complementarily, we developed the synthetic labeled reverberant LibriCHiME-5 dataset for objective evaluation with common intrusive metrics (SI-SDR, PESQ, STOI) on close-to-in-domain data. The subjective evaluation consisted of an ITU-T P.835 listening test, which was implemented in a JavaScript experimental platform and conducted in person at the University of Sheffield (UK). The subjective evaluation revealed the difficulty of the CHiME-7 UDASE task. Among the four speech enhancement systems that were evaluated in the listening test, only one succeeded in improving the overall quality compared to the unprocessed noisy speech. All systems successfully reduced the background noise but always at the expense of increased distortion, which eventually affected the overall perceived quality.
The analysis of the results revealed that the DNSMOS P.835 and TorchAudio-Squim nonintrusive performance metrics should be used with caution. Their effectiveness was demonstrated for the specific benchmarks presented in Reddy et al. [2022] and Kumar et al. [2023]. However, except for the DNSMOS BAK score, these metrics were shown to poorly correlate with the subjective ratings. On the contrary, the ranking of speech enhancement methods on reverberant LibriCHiME-5 using more traditional intrusive objective performance metrics was similar to the ranking based on subjective evaluation. This shows that the reverberant LibriCHiME-5 dataset can be used for in-domain evaluation of speech enhancement models adapted to the unlabeled CHiME-5 dataset. While this is useful practically, it is not entirely satisfying as unsupervised domain adaptation methods should in principle be evaluated on the unlabeled data of the target domain, which is very challenging methodologically [You et al., 2019, Musgrave et al., 2021]. Further research is needed to address the challenge of training and evaluation of speech enhancement models without clean speech labels.
Appendix A Experimental platform for the listening test
Screenshots of the JavaScript experimental platform developed for the listening test of the CHiME-7 UDASE task are shown in Figure 2.3.
(a) SIG rating scale and instruction.
(b) BAK rating scale and instruction.
(c) OVRL rating scale and instruction.
Figure 6: Windows of the experimental platform used for the listening test. Each window presents an instruction and a rating scale following the ITU-T P.835 recommendation, and a 5-position slider to register the vote.
Acknowledgments
The authors thank the CHiME Steering Group for their support in the organization of the CHiME-7 UDASE task and the participants of the listening test organized at the University of Sheffield. S. Leglaive acknowledges support from the French National Research Agency (ANR) as a part of the DEGREASE project (ANR-23-CE23-0009).